Hey guys, im stuck on this word problem and was wondering if someone can show me step by step on how to solve it, thanks in advance.

The Marine Corps War Memorial in Arlington Virgina, portrays the raising of the U.S flag on Iwo Jima during World War II. Find the measures of the two angles shown if the measure of <1 is 15 degrees less than twice the measure of <2

Answer by Theo(13342)

the measure of angle 1 is 15 degrees less than two times the measure of angle 2.
let angle 2 = x.
the measure of angle 1 is therefore equal to 2x - 15
presumably the 2 angles are the angles that the flag makes with the ground.
the sum of the angles is therefore equal to 180 degrees.
you get:
x + 2x - 15 = 180
simplify to get:
3x - 15 = 180
add 15 to both sides to get:
3x - 195
divide both sides by 3 to get:
x = 65
2x-15 = 115
115 + 65 = 180
all the numbers check out;.
one of the angles is 65 degrees.
the other angle is 115 degrees.
the 115 degree angle is equal to 2 times 65 - 15 = 130 - 15 = 115.
if you look at the picture in the memorial, that looks about right.
see below:
$$$